OpenStudy (anonymous):

HELP NEEDED! If f(x) = 1 over 3x − 2 and g(x) = −2x − 1, find (f • g)(9) 19 −18 0 −19 If r(x) = 3x − 1 and t(x) = −4x − 1, find (r + t)(7). 9 −9 −49 49 If f(x) = −2x − 4 and g(x) = x + 4, find (f over g)(−8). −3 3 −1 1 I don't understand how to do these in particular and it would be great if someone could help me with them.

OpenStudy (akashdeepdeb):

These questions are pretty simple [TRUST ME] When they have given functions such as f(x) = Blah blah.... g(x) = blah blah And when they tell you to calculate (f.g)(2) = Multiply both functions and put value of 2 in place of x (f/g)(3) = Divide both functions and put value of 3 in place of x (f + g)(a) = Add both functions and put value of a in place of x (f-g)(constant) = Subtract both functions and put value of the constant in place of x. Understood? :) Now do the above questions!! :D

OpenStudy (jdoe0001):

famous last words [TRUST ME]

OpenStudy (akashdeepdeb):

For eg. I'll do the first question and then you do the next 2. 1) f(x) = 1 over 3x − 2 and g(x) = −2x − 1, find (f • g)(9) (f.g)(x) = \(\frac{1}{3x-2} * {-2x-1}\) (f.g)(x) = \(\frac{-2x-1}{3x-2}\) (f.g)(9) = \(\frac{-2(9)-1}{3(9)-2}\) (f.g)(9) = \(\frac{-19}{25}\) Understood? :)

OpenStudy (jdoe0001):

keep in mind that \(\bf (r + t)(7) \implies r(7) + t(7)\quad \textit{and}\quad \cfrac{f}{g}(-8) \implies \cfrac{f(-8)}{g(-8)}\)

OpenStudy (akashdeepdeb):

[TRUST ME] LOL @jdoe0001 :D

OpenStudy (jdoe0001):

[read my lips] [I've inhaled the intoxicating fumes from the oracle of Delphi] [ so trust me]

OpenStudy (jdoe0001):

hehe

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."